If the item costs $139.99 and the item is on sale for 25% off , how much does the item cost? and how do you know ? Answer by nerdybill(7384) 
 If the item costs $139.99 and the item is on sale for 25% off , how much does the item cost? and how do you know ?
.
When you see "%" -- think "per one hundred". Therefore 25% is the same as 25/100 or .25.
.
So, the original price is $139.99
"25% off" is
25/100 * 139.99 = .25 * 139.99 = $35.00 (off)
.
"Cost" = "original cost" - "discount"
"Cost" = 139.99 - 35.00 = $104.99